# Leverage Knowledge Graph and Large Language Model for Law Article Recommendation: A Case Study of Chinese Criminal Law

### v2

@@ 4.2 CLAKG Construction and Graph Embedding
Law article id 452 Agree with 997
@@ 4.3 Case Study
4.4 Test Result
4.4.1 Analyzing the Impact of Label Imbalance in Training
We utilized the BERT[23], GRU[8] and DPCNN[11] model for law article recommendation as comparison methods. These methods were compared against the method proposed in this paper: the LLM(specifically OpenAI’s ChatGPT-4.0) for law article recommendation based on CLAKG (with case information).
Table 5: Model accuracy comparison between classical models and ours
@@ 4.4.1 Analyzing the Impact of Label Imbalance in Training
LLM + CLAKG (OURS, no case) 0.676
As shown in the Table 5, the accuracy of the BERT (baseline) is only 0.289. Upon analyzing the results provided by the BERT method, we discovered that this method classified all test cases under ”Article 133.” We believe this could be due to the small size of the training set, an imbalance in training labels[40], or an inappropriate loss function. Even after replacing the cross-entropy loss function with the Focal Loss function, the model still classified all test cases as ”Article 133”.
The GRU (baseline) model achieved an accuracy of 0.427. GRU, or Gated Recurrent Unit, is a type of recurrent neural network (RNN) that is widely used for sequential data, and this result represents its performance on the task at hand.
The DPCNN (baseline) model, which stands for Deep Pyramid Convolutional Neural Network, performed slightly better, with an accuracy of 0.495. DPCNN is a convolutional neural network architecture designed for text classification tasks, and its accuracy reflects its ability to capture hierarchical patterns in the input data.
The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than any one of the classic methods. This indicates that the proposed approach effectively mitigates the impact of the insufficient data and label imbalance in training.
4.4.2 Analyzing the Impact of Hallucinations in LLMs
We utilized the LLM[41] (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods), TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28] for law article recommendation as comparison method. This method was compared against the method proposed in this paper: the LLM for law article recommendation based on CLAKG (without case information).
Table 6: Comparing proposed method with other RAG method
@@ 4.4.2 Analyzing the Impact of Hallucinations in LLMs
As shown in the Table 6, the accuracy of the LLM without external legal data is significantly higher than that of the classic model, achieving an accuracy of 0.549. Graph-RAG’s extremely low result may be due to its graph construction or retrieval strategy not aligning well with our dataset, indicating challenges in directly applying that method to this legal domain. The proposed model (LLM+CLAKG) achieved an accuracy of 0.676, which is significantly higher than that of the TFIDF-RAG, Light-RAG, Graph-RAG method. This indicates that the proposed approach effectively mitigates the hallucinations in LLMs. We believe this is because other RAG method only matches at the text level, while the latter matches the new case with keywords at the semantic level. Additionally, the latter utilizes all information within the CLAKG through graph embedding techniques, allowing it to calculate the most relevant law articles from a holistic perspective.
@@ 4.4.3 Ablation Study
LLM + CLAKG (OURS, with case) 0.694
@@ 5 Conclusion
This paper proposes an efficient law article recommendation approach using a Case-Enhanced Law Article Knowledge Graph (CLAKG) combined with a Large Language Model (LLM). CLAKG integrates law articles and case information. It is characterized by a rich topological structure formed through usage relationships and shared key information found in judgments, which enhances the effectiveness of law article recommendation tasks. The proposed approach integrates LLMs with CLAKG, enabling more accurate recommendations of law articles and related cases by utilizing macro semantics to mitigate LLM hallucinations. The effectiveness of the proposed approach is verified through comprehensive comparisons with several baseline models on the law article recommendation task. This approach effectively addresses challenges such as insufficient data, imbalanced labels, and LLM hallucinations, leading to a significant improvement in the accuracy of the law article recommendation task.
In future work, we plan to expand the data volume of CLAKG, particularly by incorporating more case data to further enhance the performance of law article recommendation. Additionally, we will explore further applications of the CLAKG-LLM integration to improve the efficiency of court judgments.

### v3

@@ 4.2 CLAKG Construction and Graph Embedding
Law article id 452 Agree With 4820
@@ 4.3 Case Study
4.4 Experimental Results
In this subsection, we organize the baselines into three comparison groups to evaluate the proposed LLM+CLAKG framework on the law article recommendation task. Group (i) consists of classical text-classification models that directly map judgment texts to law-article labels (e.g., BERT, GRU, and DPCNN). Group (ii) includes LLM-based and generic RAG-style pipelines adapted from recent retrieval-augmented generation paradigms (e.g., plain LLM prompting, TFIDF-RAG, Graph-RAG, and Light-RAG), which represent the current mainstream approach for knowledge-intensive NLP tasks. Group (iii) presents an LLM-agnostic ablation across backbone models, where we instantiate the same CLAKG-RAG pipeline with two representative backbones—one mainstream proprietary model (GPT-4 [33]) and one strong open-source model (Llama-3.1-405B-chat [19])—and evaluate each backbone with and without CLAKG-based augmentation. This three-group organization enables a clearer positioning of our method with respect to (a) traditional neural classifiers, (b) recent generic RAG baselines, and (c) the incremental benefit of CLAKG when plugged into different LLM backbones under a unified evaluation setting.
Our primary objective here is to quantify the performance gain brought by our knowledge-graph–based RAG framework (CLAKG-RAG) for law article recommendation. In principle, any LLM can serve as a baseline for this task by directly prompting it with case facts and asking it to output applicable law articles. Given the rapid proliferation of LLM families and the non-trivial monetary and engineering cost of large-scale evaluations, a comprehensive comparison against “all” available LLMs is practically and economically infeasible. We therefore adopt a representative strategy: we keep the CLAKG construction, retrieval pipeline, and prompting scheme fixed, and instantiate them with a small but diverse set of competitive backbone models to demonstrate both effectiveness and generality. This design also reflects the model-agnostic nature of our framework: in principle, any sufficiently capable instruction-following LLM (open-source or proprietary) can be plugged into the same CLAKG-RAG pipeline in future work.
4.4.1 Comparison with classical text-classification baselines
To analyze the impact of data scarcity and label imbalance on law article recommendation, we first compare our framework with three widely used neural models for text classification, namely BERT [14], GRU [8] and DPCNN [24]. All of them treat law article recommendation as a text classification problem on the raw judgment documents. In this setting, our method corresponds to the LLM-based recommendation framework built on CLAKG with case information.
Table 6: Accuracy of classical models vs. our LLM+CLAKG.
@@ 4.4.1 Comparison with classical text-classification baselines
LLM + CLAKG (ours, with case) 0.694
As shown in Table 6, the accuracy of the BERT baseline is only 0.289. A closer inspection of the predictions reveals that BERT tends to assign almost all test cases to the majority label “Article 133”. This collapse is likely caused by the small size of the training set, the severe label imbalance, and the difficulty of capturing fine-grained legal distinctions from limited data.
The GRU baseline achieves an accuracy of 0.427. As a recurrent model designed for sequential data, GRU can capture some contextual information in the judgment texts, but it still struggles to distinguish between articles under highly imbalanced labels. The DPCNN baseline, a deep pyramid convolutional architecture tailored for text classification, further improves the accuracy to 0.495 by modeling hierarchical local patterns, yet it remains far from satisfactory. These observations indicate that, despite the moderate dataset size, the law article recommendation task on our corpus remains challenging: the label space is large (85 distinct articles in the test set) and highly imbalanced, and classical text-classification models tend to collapse to the majority label under such conditions.
In contrast, our LLM+CLAKG model (with case information) reaches an accuracy of 0.694, significantly outperforming all classical baselines. This suggests that (i) explicitly introducing legal structure via CLAKG and (ii) leveraging LLMs to reason over graph-grounded evidence can effectively mitigate the negative impact of insufficient data and label imbalance, whereas pure text-classification models are prone to majority-label bias on this dataset.
The classical text-classification models have the advantage of being conceptually simple, computationally efficient, and easy to deploy once a sufficiently large and well-balanced labeled dataset is available. However, they operate purely at the document-text level and do not encode the structured relations between cases, key legal elements, and law articles. As a result, they are highly sensitive to label imbalance and data sparsity, and tend to collapse to majority labels in our setting. In contrast, our LLM+CLAKG framework explicitly injects legal structure through the knowledge graph and leverages LLM reasoning over graph-grounded evidence. This design yields better robustness under limited and imbalanced data, at the cost of a more complex pipeline that requires KG construction and LLM inference.
4.4.2 Comparison with RAG-based baselines
To investigate the impact of hallucinations in LLMs and to compare with more recent law article recommendation paradigms, we further evaluate several RAG-based methods: a plain LLM [33] (OpenAI’s ChatGPT-4.0), TFIDF-RAG [1], Graph-RAG [16] and Light-RAG [20]. These methods instantiate retrieval-augmented generation pipelines that have recently attracted considerable attention, and here they are adapted to the law article recommendation setting. In this group of experiments, our method corresponds to the LLM-based recommendation framework built on CLAKG without using case nodes (only statutory information), denoted as LLM+CLAKG (ours, no case).
Table 7: Accuracy comparison with LLM and RAG-based baselines.
@@ 4.4.2 Comparison with RAG-based baselines
As shown in Table 7, a plain LLM without external legal data already achieves an accuracy of 0.549, which is much higher than that of the classical text-classification models. This confirms that large pretrained language models are able to capture a considerable amount of legal knowledge from their pretraining corpora. However, the LLM still suffers from hallucinations, occasionally citing inapplicable or even nonexistent articles.
TFIDF-RAG improves the accuracy to 0.597 by retrieving judgment documents and law texts based on lexical similarity and feeding them as additional context to the LLM. Light-RAG further increases the accuracy to 0.636 by using a more efficient retrieval and organization strategy. Both methods demonstrate the benefit of grounding the LLM on retrieved legal documents. In contrast, Graph-RAG performs poorly on our dataset (0.179), which we attribute to its generic graph construction and retrieval strategy not aligning well with the fine-grained structure of Chinese criminal law and the limited size of our corpus.
Our LLM+CLAKG (ours, no case) achieves an accuracy of 0.676, outperforming the plain LLM, TFIDF-RAG, Graph-RAG and Light-RAG. Compared with TFIDF-RAG and Light-RAG, which mainly rely on surface-level text similarity, our approach matches new cases to law articles via key information nodes in CLAKG, and uses graph embeddings (Section 3.2.1) to exploit the topological structure information of the graph. As a result, candidate law articles are selected based on their structured legal relevance rather than purely lexical overlap.
Moreover, by constraining the LLM to reason over candidate articles and historical cases retrieved from CLAKG (Section 3.2.4), the system is grounded in verifiable law articles and case information, which mitigates hallucinations in article citations and explanations more effectively than generic RAG pipelines. In summary, the comparison with both classical classifiers and recent RAG-based methods shows that our LLM+CLAKG framework better captures the underlying legal reasoning and achieves consistently higher recommendation accuracy.
RAG-based methods mitigate hallucinations by retrieving external texts and are relatively easy to adapt across domains. TFIDF-RAG and Light-RAG work well when lexical similarity is a good proxy for relevance, but they mainly exploit surface-level word overlap and do not explicitly model legal entities or relations; their performance may degrade when key legal distinctions are not reflected in simple bag-of-words similarity. Graph-RAG introduces a graph structure, yet its generic document graph and retrieval strategy are not tailored to the fine-grained structure of Chinese criminal law in our corpus, which limits its effectiveness. In contrast, our proposed LLM+CLAKG framework constructs a task-specific CLAKG schema centered on law article recommendation. This schema leverages legally meaningful key information in law articles together with adjudicated case information to precisely characterize the relations among different law articles as well as the links between cases and law articles. During the retrieval-augmentation stage, this semantic and structured information is injected into the LLM through prompt engineering. In this way, we substantially mitigate hallucinations of large language models and significantly improve the accuracy of law article recommendation. The trade-off is that our approach requires schema design and KG construction effort, as well as maintaining the KG over time, but it delivers consistently higher accuracy of legal reasoning than these more generic RAG pipelines.
@@ 4.4.3 Ablation Study
LLM + CLAKG (GPT-4, with case) 0.694
@@ 5 Conclusion
This paper presents an efficient law article recommendation approach that combines a Case-Enhanced Law Article Knowledge Graph (CLAKG) with a Large Language Model (LLM). CLAKG integrates law articles and adjudicated cases under a task-oriented schema, with a rich topological structure formed by usage relations and shared key information extracted from judgments. By grounding the LLM in this graph-structured macro semantics, the proposed framework produces more accurate recommendations of law articles and related cases, while mitigating typical LLM hallucinations. Extensive experiments against strong text-classification baselines (BERT, GRU, DPCNN) and recent RAG-based methods (TFIDF-RAG, Graph-RAG, Light-RAG) demonstrate that our approach effectively alleviates issues such as insufficient data, label imbalance, and hallucination in LLMs, leading to substantial gains in recommendation accuracy.
Looking ahead, we plan to expand CLAKG in both scale and scope. As more criminal judgments under Amendment (XI) become publicly available, we will be able to evaluate the framework on larger and more diverse corpora. When appropriate data access is possible, extending the evaluation to other jurisdictions will further test the scalability and generality of our approach. Beyond passively waiting for new real-world judgments, a complementary research direction is to explore carefully controlled data augmentation via large language models: LLMs can be used to generate hypothetical criminal case fact patterns conditioned on specific law articles, after which legal experts verify, correct, or discard these synthetic cases before they are integrated into CLAKG as new training and evaluation samples. Such human-validated synthetic cases have the potential to mitigate label imbalance and sparsity while preserving legal plausibility.
It is important to emphasize that the closed-loop human–machine collaboration proposed in this paper is a deployment-level mechanism for maintaining and improving CLAKG over time, rather than a separate algorithmic baseline. Accordingly, our quantitative experiments in Section 4 focus on recommendation accuracy under different model variants, assuming this closed-loop mechanism is in place. A natural next step is to design quantitative evaluation protocols for the human feedback loop itself—for example, tracking the reduction in the proportion of recommendations that require correction, measuring expert time and effort, and analyzing how the evolving CLAKG affects downstream accuracy over longer periods.
The current CLAKG construction still relies on LLM-based extraction of key information and case descriptions. Although we constrain the LLM with a predefined schema and strict post-processing, extraction errors inevitably remain and primarily manifest as missing or noisy links in the graph. Developing more robust and formally guaranteed information-extraction pipelines (e.g., combining supervised or rule-based extraction with LLMs) is therefore an important direction for future work.
While our experiments focus on Chinese criminal law, a civil-law system with codified statutes, the proposed CLAKG–LLM framework is conceptually applicable to other legal systems as well. Adapting it to different jurisdictions, however, raises several challenges. First, the underlying legal structure may differ substantially: common-law systems rely more heavily on precedents and citation networks between cases, which would require extending the CLAKG schema with additional entity and relation types (e.g., precedent cases, holdings, ratios, and explicit citation links) and redesigning task-specific key information nodes. Second, legislative drafting conventions and document structures vary across countries (e.g., hierarchical numbering, cross-references, annotations), so the extraction–transformation–loading (ETL) pipeline and schema-design steps must be tailored to each legal corpus, as illustrated by recent work on Italian legislation [38, 9]. Third, applying the framework to other languages calls for high-quality LLMs and tokenization tools for the target language, as well as careful adaptation of prompts and extraction templates to local legal terminology. Despite these challenges, the overall architecture—schema-driven KG construction, graph-embedding-based retrieval, and KG-grounded LLM reasoning—remains unchanged, suggesting that our approach can serve as a blueprint for building similar systems in other jurisdictions.
Data Availability
All source code, the constructed Case-Enhanced Law Article Knowledge Graph (CLAKG), and the processed law article and case datasets used in this paper are publicly available in a GitHub repository maintained by the authors:
The repository includes: (i) dump files of CLAKG, (ii) the de-identified structured data used in our experiments, and (iii) the full implementation of the proposed LLM–CLAKG law article recommendation framework together with instructions for reproducing the experimental results. The original raw judgments remain accessible through the official “China Judgments Online” website, subject to its access and privacy policies.